I really need help with this! I don't understand it and I have been working at it for like 30 minutes now thanks
Carlos wants to build a hot tub with a volume of 1000 cubic feet. What is a good approximation for the radius of the hot tub?

Answer by mananth(16949)

We have to assume the height.
The standard height is 3 feet.
volume of tub = pi*r^2*h
1000= pi*r^2*h
1000/(pi*h)=r^2
1000/(pi*3)=r^2
r= sqrt(1000/(3*pi))
r= 10.3 feet
..
if you increase height the diamter will decrease